# MoESD: Mixture of Experts Stable Diffusion to Mitigate Gender Bias

### v1

@@ 3.5 Mixture of Experts
MoE (Mixture of Experts) Jacobs et al. (1991); Eigen et al. (2013) is an ensemble learning technique that improves performance by weighting the predictions of different experts through gate mechanisms.
In a classic MoE system, each expert is independent, and demonstrates high performance within their area of expertise. A gating mechanism is learned to adjust the weight assigned to each expert based on the input data. In contrast, in our approach, the gates and experts, defined in Section 3.3 and 3.4, are not trained. We utilize pre-trained models, which can be flexibly replaced by other pre-trained models.
@@
To summarise, our approach is as follows: we first conduct zero-shot and unsupervised classification. We do not use the labels (the gender statistics of the 250 images for each occupation generated by the vanilla Stable Diffusion, provided by Fair Diffusion) to train the model, but only for hyperparameter selection. In other words, we only use the prompt and pretrained CLIP encoder itself. At inference, the model takes the original prompt and special token as inputs. It judges the gender skew the prompt shows, with the male skew leading to a higher probability of calling the female expert, and vice versa.
@@ 4.1 Setup
We first use the statistical data of different occupations generated by Fair Diffusion (Stable Diffusion 1.5) to select the best hyperparameters for both versions of Stable Diffusion. If the general statistical count shows more males than females, we assume that it has a male skew, and vice versa. We found that 4000 is the optimal hyperparameter $\lambda$ in Equation 1 and the subsequent derivation for version 1.5, and 100 is optimal for version 2.1. There is a large difference in the best hyperparameter between the two versions, which we attribute to the differences in the text encoder between the two versions of Stable Diffusion, which introduces different embedding gender biases in the latent space.
@@
After fine-tuning, we proceed to the inference steps: There are three experts in the system: one original, one male BiAs expert, and one female BiAs expert. When inputting a prompt, we add the same special token (“sks”) as in the fine-tuning stages. The system judges the skew, and if the skew is male, we allocate 10% to the male expert, 50% to the female expert, and 40% to the original expert. Conversely, if the skew is female, we allocate 10% to the female expert, 50% to the male expert, and 40% to the original expert. These are the optimal hyperparameters for conservative mitigation within a limited search range in our experiments since we do not want the BiAs to completely change the original weights.
In this manner, we generate 100 images based on the same prompt for each of the 153 occupations.
We then employ BLIP2 Li et al. (2023) VQA task to conduct fairness evaluation and Laion-aesthetic linear classifier Schuhmann (2022) to perform aesthetic evaluation, as we discuss in the following sections. We also employ human evaluation for the aesthetic and image-description relevance evaluation.
@@ 4.2 Fairness Evaluation
4.2.1 Metric
Stable Diffusion learns a conditional distribution $\hat{P}(X|Z=z)$, where $z$ represents the embedding of the prompt, and the biased nature of the dataset used to train the generative model and its architecture can impact the distribution $\hat{P}$. To quantify the bias in generative models, recent studies Choi et al. (2020); Teo et al. (2024); Chuang et al. (2023) propose using statistical parity. Specifically, given a classifier $h:X\rightarrow A$ for the gender attribute, the discrepancy of the generative distribution $\hat{P}$ can be defined as two metrics: the empirical and uniform distributions.
@@ 4.2.1 Metric
4.2.2 Results
We compare our method to the original Stable Diffusion, the Biased Prompt method proposed by Chuang et al. (2023), and Fair Diffusion versions 1.5 and 2.1. Due to the variability in results related to hyperparameters reported by Chuang et al. (2023), we select the best result from different hyperparameters for their method in both versions: 0.05 for version 1.5 and 500 for version 2 (as reported in their paper). The results are shown in Table 1.
@@ 4.2.2 Results
Method Fairness Score STD Fairness Score STD
@@
Bias Prompts Chuang et al. (2023) 0.279($\lambda$ = 0.05) 0.169 0.255($\lambda$ = 500) 0.143
Fair Diffusion Friedrich et al. (2023) 0.074 0.093 0.070 0.093
@@
Table 1: Fairness Evaluation of Models. For each method, we show the Fairness Score and Standard Deviation (lower values are better) for two versions of Stable Diffusion. Our MoESD-BiAs (special token) yields the best performance in all cases except for Fair Diffusion with manual guidance Friedrich et al. (2023).
MoESD-BiAs (with special token) performs better than all others except for Fair Diffusion Friedrich et al. (2023). This is expected, since we assume that guidance from manually edited prompts is SOTA in terms of fairness as it will force the model to chose the gender specified. In our method, we do not explicitly give any gender information in the prompt (e.g. male or female). We show our mitigation examples in Appendix A.1.
@@
We can observe that in most methods, the fairness score for Stable Diffusion Version 2.1 is higher than for Version 1.5, except for Fair Diffusion, which indicates that Version 1.5 performs better in terms of fairness. For other methods, there is a huge difference between different versions, but our method achieves similarly good performance in both versions.
@@
BiAs appears to perform better than full fine-tuning regardless of whether we use the special token, which suggests that only a small set of parameters (5%) are enough to mitigate bias. We attribute it to the fact that adapters allow more targeted parameter updates, reducing the risk of gradient vanishing or exploding and preventing the catastrophic forgetting of general knowledge from well-pretrained Stable Diffusion weights. Moreover, due to random initialization, our adapters are more effective in mitigating bias in relatively small data and a few fine-tuning steps.
@@
The special token contributes substantially to mitigating bias since it helps the model (bias experts) to remember the fine-tuning process, which is especially effective for the randomly initialized adapters. We can see the MoESD method performs worse than the vanilla method. We attribute it to the fact that small fine-tuning cannot impact biases of a large base model. However, the special token can mitigate this, as seen from the MoESD (special token) results.
@@ 4.3 Aesthetic Evaluation
4.3.1 Metric
It should be noted that due to differences in training data, the style and quality of generated images vary. While the images generated by our model achieve better fairness scores than those of others, we also need to evaluate whether the aesthetic quality of the generated images is not worse.
Linear Aesthetic Evaluation
Simulacra Aesthetic Captions - SAC Pressman et al. (2022) is a dataset consisting of over 238,000 synthetic images generated with AI models, from over forty thousand user submitted prompts. Users rate the images on their aesthetic value from 1 to 10, when they were asked “How much do you like this image on a scale from 1 to 10?”. LAION-Aesthetics Schuhmann (2022) trains a linear model on 5000 image-rating pairs from the SAC dataset, which can predict a numeric aesthetic score in 1-10. We utilize this linear model to evaluate the image sets of all methods and compare their aesthetic scores.
Human Evaluation
We randomly select 30 out of 153 occupations and then randomly select one out of 100 generated images. We choose the images from Fair Diffusion Friedrich et al. (2023), Bias Prompts ($\lambda$ = 0.05) Chuang et al. (2023), and our best method (MoESD-BiAs (special token)), and compare them with the vanilla images, all implemented on Stable Diffusion Version 1.5. So there are 90 pairs of images (30 pairs compare Fair Diffusion with vanilla, 30 pairs compare Bias Prompts ($\lambda$ = 0.05) with vanilla, and 30 pairs compare ours with vanilla) for users to choose which one is better. The instruction is “Choose the image that has better quality (with ‘same’ being an option), where the main criterion is that the person in the image looks like a normal person and their there are no unrealistic items in the background”. More details of the survey are listed in Appendix A.10.
It’s important to note that in some pairs, the vanilla model and ours have the same image because there is a probability that we do not activate the Bias expert in our model. In those cases, the quality remains the same. The same happens in some instances of Fair Diffusion, where the gender prompt engineering may not lead to different results. For the Bias Prompts, the best performance of Stable Diffusion Version 1.5 occurs with a small $\lambda$ value, so the pairs should be generally similar.
As final result, we calculate the average ratio at which users think each model does not perform worse than the vanilla model.
4.3.2 Results
Linear Aesthetic Evaluation
@@ 4.3.2 Results
Method Aesthetic Score Aesthetic Score
@@
Bias Prompts Chuang et al. (2023) 5.132($\lambda$ = 0.05) 5.285($\lambda$ = 500)
Fair Diffusion Friedrich et al. (2023) 5.154 5.318
@@
Table 2: Aesthetic Evaluation of Models. For each method, we show the Aesthetic Score (higher values are better) for two versions of Stable Diffusion. Our method yields the best performance in both versions.
Compared to the vanilla approach, MoESD-BiAs (special token) has a higher aesthetics score, even though it may blurs faces at times. We speculate that the higher aesthetics score may be because of images with more colorful and complex backgrounds, which are favored by the aesthetic classifier. Additionally, we also observed that the vanilla method sometimes produces images with words, which receive very aesthetics scores, as also noted in the Laion-aesthetic example [^4] http://captions.christoph-schuhmann.de/aesthetic_viz_laion_sac+logos+ava1-l14-linearMSE-en-2.37B.html. Finally, the score for version 2.1 is higher than for version 1.5, indicating that version 2.1 has better aesthetic recognition during pre-training.
@@
Human Evaluation
We present our human evaluation results in Table 3. Our method was same or better than the vanilla method in 59% of the cases, which ranks it worse than the SOTA methods we evaluated. This reflects the fact that our images are sometimes blurry, which is a very visible artifact and leads our human evaluators to (rightly) prefer vanilla images in many cases. This lower quality may be due to the limited size of the training data and the small number of training parameters. This indicates that more work is needed to keep the quality of the generated images, while mitigating biases.
Method Non-worse Ratio
Fair Diffusion Friedrich et al. (2023) 59.3%
Bias Prompts($\lambda$ = 0.05) Chuang et al. (2023) 85.16%
MoESD-BiAs (special token) 54.58%
Table 3: Average Non-Worse Ratio for 3 Methods.
@@ 4.4 Image-Description Relevance Evaluation
We then calculated the mean and median relevance scores for each image and conducted an analysis for all 30 images. Details of the survey are listed in Appendix A.10.
The results are shown in Table 4. We observe a small degradation in quality when our bias mitigation method is activated. However, this degradation is acceptable, as the score difference is within a range of 1 (in 10).
Method Mean Median
@@
Table 4: Mean and Median of Human Evaluation We evaluate how well the image matches the image description. The number in parentheses represents the number of images.
@@ 5 Conclusions
In this paper, we introduce the MoE to Stable Diffusion by taking a major step by assessing existing biases in the prompt embedding and fine-tuning the BiAs with special tokens to aid in understanding the biased data, which results in our MoESD-BiAs system, which has achieved significant success in mitigating bias while maintaining quality.
6 Limitations
Although our method has made significant strides in mitigating bias while maintaining aesthetic quality, we must admit the limitations of our work.
First, we cannot achieve the same level of fairness score as manual editing of the gender attributes, which could be further improved. Additionally, we did not pay attention to evaluating biases related to the race, as measuring bias in prompt embeddings for the race is much more difficult than for gender. Moreover, in our work, we only address the case of binary gender. We did not consider sexual minorities in various contexts, which is a much more complex task. We hope to explore this further in future work.
Second, our zero-shot and unsupervised prompt bias identification and hyperparameter selection are based on the results of Fair Diffusion statistical counts Friedrich et al. (2023), so it may not be entirely accurate. Moreover, we only achieve 79% accuracy in identifying bias from the prompt, which is not perfect when facing challenging occupations such as the insurance agent.
Third, although we achieve good performance in aesthetics, our method indeed loses some fine-grained details on faces due to the quality of the fine-tuning images. However, fortunately, our experts can be switched by fine-tuned ones by users themselves, which can be further improved through special fine-tuning.
7 Ethical Considerations
However, by introducing the BiAs approach to mitigate bias, there is a risk that people might misuse the weights to generate more biased content.
Acknowledgements
We would like to express gratitude to all who participated in the human evaluation, which has significantly enhanced the quality of this research.
Appendix A Appendix
A.1 Mitigation Visualization
@@
A.2 Text-Encoded Bias from Prompts
@@
Figure 4: T-Distributed Stochastic Neighbor Embedding (t-SNE) Van der Maaten and Hinton (2008) dimension reduction and visualization of prompts encoded by text encoder of Stable Diffusion Version 1.5 and 2.1 Rombach et al. (2022). There is a clear boundary between two gender-bias occupation embeddings in both versions of Stable Diffusion.
A.3 Different Methods for Bias-Identification Gate
@@
We also utilize third-party models (T5 Raffel et al. (2020) and Sequence Transformer Reimers and Gurevych (2019)) as monitoring models to utilize their embedding similarity for classification and compare the results with our method.
@@
A.4 More Details of Chuang et al. (2023)
@@
$P^{*}=P_{0}\left(I+\frac{\lambda}{\left|S\right|}\sum_{(i,j)\in S}(z_{i}-z_{j})(z_{i}-z_{j})^{T}\right)^{-1}$ (8)
@@
A.5 Parameter $\lambda$ for Bias-Identification

### v2

@@ 3.5 Mixture of Experts
MoE (Mixture of Experts) (Jacobs et al. 1991; Eigen, Ranzato, and Sutskever 2013) is an ensemble learning technique that improves performance by weighting the predictions of different experts through gate mechanisms.
In a classic MoE system, each expert is independent and demonstrates high performance within their area of expertise. A gating mechanism is learned to adjust the weight assigned to each expert based on the input data. In contrast, in our approach, the gates and experts, defined in Section 3.3 and 3.4, are not trained. We utilize pre-trained models, which can be flexibly replaced by other pre-trained models.
@@
To summarise, our approach is as follows: we first conduct zero-shot and unsupervised classification. We do not use the labels (the gender statistics of the 250 images for each occupation generated by the vanilla Stable Diffusion, provided by Fair Diffusion) to train the model, but only for hyperparameter selection. In other words, we only use the prompt and pre-trained CLIP encoder itself. At inference, the model takes the original prompt and special token as inputs. It judges the gender skew the prompt shows, with the male skew leading to a higher probability of calling the female expert, and vice versa. We show our mitigation examples in Appendix A.
@@ 4.1 Setup
We first use the statistical data of different occupations generated by Fair Diffusion (Stable Diffusion version 1.5) to select the best hyperparameters for both versions of Stable Diffusion. If the general statistical count shows more males than females, we assume that it has a male skew, and vice versa. We found that 4000 is the optimal hyperparameter $\lambda$ in Equation 1 and the subsequent derivation for version 1.5, and 100 is optimal for version 2.1. There is a large difference in the best hyperparameter between the two versions, which we attribute to the differences in the text encoder between the two versions of Stable Diffusion, which introduces different embedding gender biases in the latent space.
@@
After fine-tuning, we proceed to the inference steps: There are three experts in the system: one original, one male BiAs expert, and one female BiAs expert. When inputting a prompt, we add the same special token (“sks”) as in the fine-tuning stages. The system judges the skew, and if the skew is male, we allocate 10% to the male expert, 50% to the female expert, and 40% to the original expert. Conversely, if the skew is female, we allocate 10% to the female expert, 50% to the male expert, and 40% to the original expert. These are the optimal hyperparameters for conservative mitigation within a limited search range in our experiments. We do not want the BiAs to completely change the original weights, as our Bias Identification Gate mechanism is not absolutely accurate. In this manner, we generate 100 images based on the same prompt for each of the 153 occupations.
We then employ BLIP2 (Li et al. 2023) VQA task to conduct fairness evaluation, evaluation method of Esposito et al. (2023), and Laion-aesthetic linear classifier (Schuhmann 2022) to perform aesthetic evaluation, as we discuss in the following sections. We also employ human evaluation for image-description relevance evaluation.
In our method, we do not explicitly provide any gender information in the prompt, token, or gender guidance in the latent space. Therefore, we do not compare our method with those that include gender hints and guidance.
@@ 4.2 Fairness Evaluation
Metric
Stable Diffusion learns a conditional distribution $\hat{P}(X|Z=z)$, where $z$ represents the embedding of the prompt, and the biased nature of the dataset used to train the generative model and its architecture can impact the distribution $\hat{P}$. To quantify the bias in generative models, recent studies (Choi et al. 2020; Teo, Abdollahzadeh, and Cheung 2024; Chuang et al. 2023) propose using statistical parity. Specifically, given a classifier $h:X\rightarrow A$ for the gender attribute, the discrepancy of the generative distribution $\hat{P}$ can be defined as two metrics: the empirical and uniform distributions.
@@
Results
We compare our method to the original Stable Diffusion, DebiasVL, Ethical Intervention, and Fairness Fine-tuning across versions 1.5 and 2.1. Due to the variability in results related to hyperparameters reported by DebiasVL, we select the best results from different hyperparameters for their method in both versions: 0.05 for version 1.5 and 500 for version 2.1, as reported in their paper. For Ethical Intervention, we use the prompt intervention: “if all individuals can be a [occupation] irrespective of their gender.” For Fairness Fine-tuning, we use the provided pre-trained weights of version 1.5 and report three different fine-tuning methods: fine-tuning the text encoder (TE), fine-tuning the U-Net, and fine-tuning both of them. MoESD-BiAs (with special token) performs better than all others. The results are shown in Table 1. Below we summarise our main observations:
@@ Results
Method Fairness Score $\downarrow$ STD $\downarrow$ Fairness Score $\downarrow$ STD $\downarrow$
@@
DebiasVL 0.279 0.169 0.255 0.143
Ethical Int. 0.379 0.120 0.374 0.140
Fairness Fine-tuning (TE) 0.239 0.132 - -
Fairness Fine-tuning (U-Net) 0.255 0.149 - -
Fairness Fine-tuning (TE+U-Net) 0.273 0.157 - -
@@
Table 1: Fairness Evaluation - Comparison with Vanilla Stable Diffusion (Rombach et al. 2022), DebiasVL (Chuang et al. 2023), Ethical Intervention (Bansal et al. 2022), and Fairness Fine-tuning (Shen et al. 2024). For each method, we show the Fairness Score and Standard Deviation for two versions of Stable Diffusion. Our MoESD-BiAs (special token) yields the best performance in all cases.
Occupation Robustness
Ethical Intervention performs worse than the vanilla model. This is because the method is effective in professions that are not as challenging as those listed in the profession list of Fair Diffusion. Moreover, interventions without gender attributes worsen the fairness results, indicating that excessive intervention makes it difficult for the model to handle challenging occupations. Our method, through effective fine-tuning on the more difficult occupations, demonstrates robustness in these cases.
@@
In most methods, the fairness score for Stable Diffusion Version 2.1 is higher than for Version 1.5, except for Ethical Intervention, which indicates that Version 1.5 performs better in terms of fairness. For other methods, there is a significant difference between different versions. However, our MoESD-BiAs (special token) achieves similarly good performance in both versions.
@@
In our method, BiAs appears to perform better than full fine-tuning regardless of whether we use the special token, which suggests that only a small set of parameters (5%) are enough to mitigate bias. Additionally, the results of Fairness Fine-tuning show that Fairness Fine-tuning (TE) and Fairness Fine-tuning (U-Net) perform better than Fairness Fine-tuning (TE+U-Net), indicating that excessive fine-tuning is not always beneficial in mitigating gender bias. We attribute it to the fact that adapters allow more targeted parameter updates, reducing the risk of gradient vanishing or exploding and preventing the catastrophic forgetting of general knowledge from well pre-trained Stable Diffusion weights. Moreover, due to random initialization, our adapters are more effective in mitigating bias in relatively small data and a few fine-tuning steps.
@@
The arbitrary special token contributes substantially to mitigating bias since it helps the model (bias experts) to remember the fine-tuning process, which is especially effective for the randomly initialized adapters. We can see the MoESD method performs worse than the vanilla method. We attribute it to the fact that small fine-tuning cannot impact biases of a large base model. However, the special token can mitigate this, as seen from the MoESD (special token) results.
@@ 4.3 Aesthetic Evaluation
Metric
It should be noted that due to differences in training data, the style and quality of generated images vary. Since our training dataset is AI-generated instead of using real images, we noticed that some of our images contain blurred faces. Therefore, we need to find a way to evaluate the image quality. However, due to the synthetic nature of the training dataset and generated content, we cannot use the Fréchet Inception Distance (FID) (Heusel et al. 2017) for evaluation.
Simulacra Aesthetic Captions - SAC (Pressman, Crowson, and Contributors 2022) is a dataset consisting of over 238,000 synthetic images generated with AI models, from over forty thousand user submitted prompts. Users rate the images on their aesthetic value from 1 to 10, when they were asked “How much do you like this image on a scale from 1 to 10?”. LAION-Aesthetics (Schuhmann 2022) trains a linear model on 5000 image-rating pairs from the SAC dataset, which can predict a numeric aesthetic score in 1-10. We utilize this linear model to evaluate the image sets of all methods and compare their aesthetic scores.
Results
@@
Method Aesthetic Score $\uparrow$ Aesthetic Score $\uparrow$
@@
DebiasVL 5.132 5.285
Ethical Int. 5.134 5.286
Fairness Fine-tuning (TE) 5.022 -
Fairness Fine-tuning (U-Net) 5.024 -
Fairness Fine-tuning (TE+U-Net) 5.067 -
@@
Table 2: Aesthetic Evaluation - Comparison with Vanilla Stable Diffusion (Rombach et al. 2022) , DebiasVL (Chuang et al. 2023), Ethical Intervention (Bansal et al. 2022), and Fairness Fine-tuning (Shen et al. 2024). For each method, we show the Aesthetic Score for two versions of Stable Diffusion. Our method yields the best performance in both versions.
Compared to the vanilla approach, MoESD-BiAs (special token) has a higher aesthetics score, even though it may blurs faces at times. We speculate that the higher aesthetics score may be because of images with more colorful and complex backgrounds, which are favored by the aesthetic classifier. Additionally, we also observed that the vanilla method sometimes produces images with words, which receive very low aesthetics scores, as also noted in the Laion-aesthetic example [^4] http://captions.christoph-schuhmann.de/aesthetic˙viz˙laion˙sac+logos+ava1-l14-linearMSE-en-2.37B.html. Finally, the score for version 2.1 is higher than for version 1.5, indicating that version 2.1 has better aesthetic recognition during pre-training.
@@ 4.4 Image-Description Relevance Evaluation
We then calculated the mean and median relevance scores for each image and conducted an analysis for all 30 images. Details of the survey are listed in Appendix J.
In this sanity check, we only aim to prove that our method does not significantly affect the image quality compared to the original. Therefore, we did not include other methods in the human evaluation. The results are shown in Table 3. We observe a small degradation in quality when our bias mitigation method is activated. However, this degradation is acceptable, as the score difference is within a range of 1 (in 10).
Method Mean $\uparrow$ Median $\uparrow$
@@
Table 3: Mean and Median of Human Evaluation. The number in parentheses represents the number of images.
4.5 Debiasing for Skin Tone
Here we applied our method for a different bias mitigation task: bias of race, as defined by skin color. We evaluate our method on the skin tone task of Esposito et al. (2023), which is simpler than using complex race categories. The task is similar to debiasing for gender: it uses the same profession prompts to generate images, but evaluates the images at the skin tone level. Evaluation details can be found in the Appendix H. Our results indicate that while the performance of the bias identification gate does not match the accuracy levels seen in other common binary classification tasks, it still plays a crucial role in mitigating gender and skin tone bias. We show our mitigation result in Table 4 compared with the Vanilla method in Stable Diffusion v1.5.
Method Fairness Score $\downarrow$ STD $\downarrow$
Vanilla 0.209 0.134
MoESD-BiAs (special token) 0.127 0.106
Table 4: Effectiveness of our Method on Skin Tone.
@@ 5 Conclusions
We introduced the MoESD-BiAs approach to text-to-image models: an MoE-based method integrated to Stable Diffusion to assess biases in the prompt embeddings, combined with fine-tuning the BiAs approach with arbitrary tokens to mitigate biases in the model. By adjusting the model weights to counteract inherent biases, we generate gender-fair images without requiring specific gender attribute guidance from the prompt, allowing the model to further improve when combined with other methods.
Appendix A Mitigation Visualization
@@
Appendix B Text-Encoded Bias from Prompts
@@
Figure 4: T-Distributed Stochastic Neighbor Embedding (T-SNE) dimension reduction and visualization of prompts encoded by text encoder of Stable Diffusion Version 1.5 and 2.1 (Rombach et al. 2022). There is a clear boundary between two gender-bias occupation embeddings in both versions of Stable Diffusion.
Appendix C Different Methods for Bias-Identification Gate
@@
We also utilize third-party models (T5 and Sequence Transformer ) as monitoring models to utilize their embedding similarity for classification and compare the results with our method.
@@
Appendix D More Details of DebiasVL
@@
$P=P_{0}\left(I+\frac{\lambda}{\left|S\right|}\sum_{(i,j)\in S}(z_{i}-z_{j})(z_{i}-z_{j})^{T}\right)^{-1}$ (8)
@@
Appendix E Parameter $\lambda$ for Bias-Identification